OpenStudy (anonymous):

a population of grasshoppers quadruples in 20 days. assuming exponential growth, if the present population is 40 million, what will it be in 50 days. answer by first finding the number of ghoppers as a function of time. y=y0e^kt.

OpenStudy (anonymous):

I devided 50 by 20 and got 2.5 than I multiplied 40 by 4 because I am qudrupaling it 40 times 4+40 times 4+40 times 2: this is because you quadrupal it 2.5 times so it would be: 160+160+80= 400 400 million grasshoppers

OpenStudy (anonymous):

hope this helps:D

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome:D

OpenStudy (anonymous):

lets try this: \[y=y_0e^{kt} \] \[y_0=40\]so you have \[y=40e^{kt}\]

OpenStudy (anonymous):

you do not know k, but you know that when t = 20, y=160 (quadruples) so we write \[160=40e^{20k}\] and solve for k

OpenStudy (anonymous):

first step divide by 40, it was unimportant in any case \[4=e^{20k}\] \[ln(4)=20k\] \[k=\frac{ln(4)}{20}=.0693\]rounded

OpenStudy (anonymous):

i love how helpful you are thank you so much

OpenStudy (anonymous):

now we go back to the formula and have \[y=40e^{.0693t}\] replace t by 50 to get \[y=40e^{.0693\times 50}=40e^{3.46574}=40\times 32=1280\]\]

OpenStudy (anonymous):

my answer may be off so let me check it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i made a calculation error somewhere but it is certainly not the answer you were given earlier. doubles every twenty days and when you start counting you have 40 so in 20 days you have 80 and then in another 20 days (40 days later) you have 160

OpenStudy (anonymous):

oh wait. quadruples! ok you start counting you have 40

OpenStudy (anonymous):

yeah i got 160 when i tried on my own, i just got confused

OpenStudy (anonymous):

yeah i got 160 when i tried on my own, i just got confused

OpenStudy (anonymous):

yeah i got 160 when i tried on my own, i just got confused

OpenStudy (anonymous):

in twenty day s you have 160 in another 20 day you have 640

OpenStudy (anonymous):

so where do i plug that in?

OpenStudy (anonymous):

so where do i plug that in?

OpenStudy (anonymous):

so where do i plug that in?

OpenStudy (anonymous):

ok i was right.

OpenStudy (anonymous):

everything i wrote was correct.

OpenStudy (anonymous):

we can do it the easy way without solving like we did, although that is what they want you to do

OpenStudy (anonymous):

first find k, and then replace t by 50

OpenStudy (anonymous):

it is all correct. but lets do it the easy way without all that. you know it quadruples every 20 days. so another formula you can use is \[40\times 4^{\frac{t}{20}}\]

OpenStudy (anonymous):

now just replace t by 50 and you get \[40\times 4^{\frac{50}{20}}=40\times 4^{\frac{5}{2}}\]

OpenStudy (anonymous):

\[4^{frac{5}{2}}=\sqrt{4^5}=2^5=32\]

OpenStudy (anonymous):

and \[32\times= 1280\]

OpenStudy (anonymous):

i meant 32*40=1280

OpenStudy (anonymous):

either way that is the correct answer

OpenStudy (anonymous):

okay makes sense! thanks again

OpenStudy (anonymous):

welcome